OpenStudy (anonymous):

solve \[ n-5\ge3\] \[A.) n \ge-8\] \[B.) n \ge8 \] \[C.) n \ge2 \] \[D.) n \ge-2 \]

jimthompson5910 (jim_thompson5910):

Add 5 to both sides to isolate n (you want to undo the subtraction of 5) \[\large n-5\ge 3\] \[\large n-5+5\ge 3+5\] \[\large n+0 \ge 8\] \[\large n \ge 8\]

OpenStudy (anonymous):

wow when you put it like that its easy ! thank you

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

\[14+x \le3 \] how do i do this one

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

14+x is the same as x+14

jimthompson5910 (jim_thompson5910):

how do you undo that +14

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

\[x \le-11\]

jimthompson5910 (jim_thompson5910):

good, you subtract 14 from both sides

OpenStudy (anonymous):

\[\frac{ x }{ 21}\le-6\]

OpenStudy (anonymous):

how do i do that ?

OpenStudy (anonymous):

@jim_thompson5910